OpenStudy (anonymous):

A farmer wants to enclose 135000 square yards of land in a rectangular plot. The material used for the front side costs $2 per yard, the material for the other three sides costs $1 per yard. Determine the dimensions that will minimize cost.

OpenStudy (anonymous):

ok i have gotten as far as: 3-(270000/y^2) = 0 do I solve for y?

OpenStudy (anonymous):

call the side that cost $1 , and the side that is $2 y, then you have a total cost is \(C=2x+3y\) and you also know that \(xy=13500\) and therefore \(y=\frac{135000}{x}\) replace in first equation to get \[C(x)=2x+\frac{3\times 135000}{x}\]

OpenStudy (anonymous):

ok i screwed that up

OpenStudy (anonymous):

should be \[C=3x+2y\] \[C(x)=3x+\frac{2\times 135000}{x}\]

OpenStudy (anonymous):

oh i see you have gotten past this step and have taken the derivative and now just need to solve

OpenStudy (anonymous):

\[3-\frac{270000}{x^2}=0\] \[\frac{270000}{x^2}=3\] \[270000=3x^2\] \[x^2=90000\]

OpenStudy (anonymous):

so \(x=300\)

OpenStudy (anonymous):

x=3000

OpenStudy (anonymous):

yes thats where i was confused satellite73

OpenStudy (anonymous):

i wasnt solving for y correctly

OpenStudy (anonymous):

y=45